# Parameter-free prediction of irradiation defect structures in tungsten at room temperature using stochastic cluster dynamics Sicong He1, Brandon Schwendeman2, George Tynan2, Jaime Marian1,3

### v1

@@ 2.3.2 Cluster size distribution
While the total number of Frenkel pairs per PKA is the starting point of the defect insertion process, another key piece of information concerns the sizes of the defect clusters inserted in the simulation volume with every displacement cascade. Detailed studies based on MD simulations have suggested that cluster sizes follow a universal scaling law independent of PKA energy [56, 26]. This is substantiated by recent results by Byggmastar et al. [54] including cascades with PKA energies up to 2 MeV. The normalized histograms based on the full dataset are shown in Figure 4(a) for vacancy, Fig. LABEL:vac-dist, and SIA, Fig. LABEL:sia-dist, clusters. We have fitted the distributions to inverse power laws with scaling , where is the number of point defects in the cluster and is the scaling exponent. As the plots show, the exponents are and for vacancy and SIA clusters, denoted as and , respectively. We use these values in our simulations to allocate the number Frenkel pairs generated by each PKA using the damage correlations in eqs. (2) to defect clusters, until the total number of defects per cascade is fully apportioned.
@@ 2.3.3 Mobile defect species
As discussed above, the majority of defect clusters larger than a given size adopt a prismatic loop configuration. These loops are defined unequivocally by their Burgers vector and by their habit plane [57, 58]. While in principle all perfect (unfaulted) loops are glissile along their Burgers vector’s direction, in W they are practically immobile for clusters with more than three interstitials or one vacancy [59]. Indeed, there is consistent evidence from isochronal annealing studies that the only species contributing to recovery in stages I, II, and III are, respectively, single SIA, di-interstitials, and monovacancies [45, 46, 60, 43, 47, 48, 44]. TEM experiments in high-purity electron irradiated W have shown that clusters move only in highly driven conditions, distinct from conventional ion beam irradiation conditions [61]. Even in atomistic simulations, in pure lattices with no imupirities or imperfections, it has been shown that small SIA clusters have a reduced mobility compared to other BCC metals [62]. Object kinetic Monte Carlo simulations also point to limited or no SIA cluster mobility in W to find consistency with experimental defect concentration measurements [24]
Despite some contradicting work on the diffusion properties of single self-interstitials, in which migration energies as high as 0.3 eV have been proposed [15, 63, 64, 65], we have found that the migration energies of SIA that lead to agreement with the present experiments (and are consistent with additional experimental measurements [66] should be 0.01 eV. Indeed, one of the main reasons for these low SIA migration energies is that their migration is almost exclusively one-dimensional along directions [66]. It is known that considering rotation as part of the transport mechanism of SIA leads to larger effective migration energies, as several atomistic calculations have demonstrated [67, 15, 64, 68]. However, this is at odds with experimental observations [69, 66, 70] (and other conflicting atomistic studies [67, 14, 59]). As well, some authors have suggested that the reason for this low migration energies is that quantum effects may suppress the temperature dependence at low temperatures in W [71]. With that in mind, the diffusivity parameters for the mobile species in our model, and the source for their values, are as listed in Table 2.
@@
| 0.009 | [59, 70] | ||
| 0.024 | [59] | ||
| 1.71 | [72, 17, 73, 70, 74] |
2.4 One-dimensional migration of single self-interstitials
As indicated in the previous subsection, SIA mobility is a key property governing the agreement between the experimental measurements and the modeling predictions. With the migration properties firmly established, here we characterize its other crucial aspect: 1D diffusion333Note that what is referred to as ‘1D diffusion’ in the context of irradiation damage is the migration along rectilinear paths of defects evolving in a 3D setting.. Consideration of one-dimensional migration of self-interstitials implies employing material parameters in eq. (5) that reflect this characteristic of SIA motion. In particular, 1D migration leads to specific coefficients and . In our setting, where the only operational defect sinks are free surfaces bounding the irradiated sample film, the corresponding sink strength, , for SIA migration is [75, 76]:
where is the thickness of the irradiated film. The reaction rate between a 1D-moving species (e.g., a single-interstitial) and an immobile object (a vacancy, vacancy cluster, or another SIA cluster) is written as [76]:
where is the diffusivity of (mobile) species , and are the interaction radii of the two reacting species, and and are their respective concentrations. Note that, for low values of , the above rate closely resembles that of the 3D-3D case. After Huang and Marian [76], due to their extremely low interaction cross-section, the reaction rate between two one-dimensionally migrating species reactions (e.g., single SIA reacting with one another) is neglected in this work.
@@ 2.4.1 PKA energy distribution tail
The final element in our model that has a strong bearing on the ability of our simulations to capture the experimental size distributions is the need to introduce PKA with sufficiently-high energies to sample the large-size portion of the cluster size distribution. As justified above, due to the limited defect mobility at 290 K in W, the only feasible formation mechanism of large loops in this case is by direct injection in displacement cascades.
However, as noted by Sand et al. [56] and Byggmastar et al. [54], the probability of generating loops with more than 100 defects in W requires PKA energies higher than 100 keV. This is consistent with the expressions used in this work, given in eq. (2). From Fig. 2, the probability of sampling energies larger than 100 keV is one in 100,000. Our SRIM simulations indicate that each ion on average creates recoils, i.e., a minimum of six ion insertions is needed to produce at least one 100-keV PKA. The ion insertion rate, , is calculated from the parameters in Table 1 as:
where is the dpa rate, the atomic density of W, and the simulation volume referred to in A. For sufficient statistics, it is desirable to accelerate the ion insertion rate as much as possible, which implies using as large a value of as reasonably possible. In this work we use cm-3.
@@ 3.1 Cluster size distributions
Figure 6(a) shows the computed SIA and vacancy cluster size distributions at the end of the irradiation period and after a room temperature annealing of one hour of duration. In both cases, the experimentally measured defect size distributions (corresponding to the ‘after annealing’ simulation results) are shown for comparison.
Figure LABEL:sia-scatter displays the SIA cluster size distribution in the ‘as irradiated’ and ‘after annealed’ conditions. The scatter in the model defect size distributions is due to the stochastic nature of our simulations, which, for a given parameter set, yield different but statistically valid results when repeated several times. Furthermore, the statistical variance increases with cluster size due to limited statistics. As the figure shows, the difference between the as-irradiated and annealed size distributions is most visible for cluster sizes with . This is a direct consequence of the high mobility of and defects, whose populations, in the absence of irradiation, are partitioned in the following manner:
@@
Our results show that the defect size distributions match the experimental distribution for large cluster sizes. Interestingly, the agreement is remarkably accurate if one considers the lower edge (convex hull) of the simulated statistical distributions. For smaller defects, the agreement improves after the simulated annealing, owing principally to a decline in the populations of small SIA clusters through dissociation and recombination.
Figure LABEL:vac-scatter shows the vacancy cluster size distribution functions for the as-irradiated and annealed conditions. The figure shows the extent of TEM size resolution regions assuming planar loos () (as a reference, for globular clusters, ). Qualitatively, the vacancy cluster population displays a size distribution similar to its self-interstitial counterpart, with large scatters in the cluster populations for large defect sizes. However, in this case, unlike for SIA defects, the plot shows a stable monovacancy contribution in the as-irradiated condition that reaches very high concentrations. This is a consequence of the negligible mobility of vacancies at 290 K, which leads to a high retention of monovacancies in the material. After the annealing, most of the di-vacancy () population disappears on account of its thermal instability (negative binding energy, see Table 3), and a concentration maximum develops for the cluster. For its part, the experimental size distribution underestimates the concentrations of vacancy clusters with , both in the as-irradiated and annealed conditions, but displays excellent agreement with the simulation results at larger cluster sizes.
@@ 3.2 Defect cluster evolution
Figure 6(b) shows the evolution of the SIA and vacancy subpopulations as a function of irradiation time. Each population is partitioned into single point defects ( and ), all clusters (, ), and loops and voids above the TEM resolution limit ( and , respectively). The evolution of SIA defects is displayed in Fig. LABEL:sia-concentration, where the unstable nature of the single SIA population is clear from its relatively low concentrations and highly volatile time evolution. This is, as discussed above, a direct consequence of their high mobility and transient behavior. In contrast, the SIA cluster populations exhibit well-defined, monotonically-increasing trends throughout the entire irradiation process.
The evolution of the vacancy clusters is shown in Fig. LABEL:vac-concentration. In this case, the monovacancy population is seen to be stable and grow at the same rate as larger clusters, although in much larger concentrations. Indeed, more than 90 Clusters of both types are seen to increase linearly, indicating that the system is found in a transient stage of defect accumulation and has not reached saturation yet.
@@ 4.1 Capabilities of physics-based numerical simulation model
We have identified four essential factors without which model predictions would not have achieved the high degree of agreement with the experimental data. In no particular order, these factors are: (i) accurately capturing the amount of defects produced in cascades as a function of PKA energy; (ii) one-dimensional, athermal SIA diffusion; (iii) universal scaling law of cluster sizes for defects generated in displacement cascades (independent of PKA energy); and (iv) consideration of the disproportionate importance of the high-energy tail of the PKA energy distribution function. All four have been implemented in our models, jointly contributing to the accuracy of the simulations and the agreement with experiments. Point (i) ensures that a sufficient number of defects are inserted to account for the large clusters observed experimentally. Point (ii) maintains a high SI mobility at low temperatures, which controls recombination with vacancies and post-cascade SIA loop growth. Point (iii) guarantees that the defect clusters inserted into the simulation volume appear with the correct sizes from the beginning. Finally, point (iv) plays the crucial role of ensuring that the material is seeded with few but important large immobile dislocation loops, which act as a stable substrate on which further irradiation damage accumulates and evolves.
The four factors above pertain to the physics of irradiation damage in W, and are in principle agnostic to the model through which they are implemented. These may include okMC [14, 4, 29], MFRT [78, 24, 79], or our hybrid method SCD, which builds on advantageous features of both: (i) a mean-field description of the defect populations for computational efficiency and reaching high irradiation doses, and (ii) uses a kinetic Monte Carlo solver to avoid numerical solution bottlenecks in MFRT and provide a meaningful statistical variance to assess numerical errors.

### v3

@@ 2.3.2 Cluster size distribution
While the total number of Frenkel pairs per PKA is the starting point of the defect insertion process, another key piece of information concerns the sizes of the defect clusters inserted in the simulation volume by each displacement cascade. Detailed studies based on MD simulations have suggested that cluster sizes follow a universal scaling law independent of PKA energy sand2013high, mason2019atomistic. This is supported by recent results by Byggmastar et al. byggmastar2025four including cascades with PKA energies up to 2 MeV. The normalized histograms based on the full dataset are shown in Figure 4(a) for vacancy, Fig. LABEL:vac-dist, and SIA, Fig. LABEL:sia-dist, clusters. We have fitted the distributions to inverse power laws with scaling , where is the number of point defects in the cluster and is the scaling exponent. As the plots show, the exponents are and for vacancy and SIA clusters, denoted as and , respectively. We use these values in our simulations to allocate the number Frenkel pairs generated by each PKA using the damage correlations in eqs. (2) to defect clusters, until the total number of defects per cascade is fully apportioned.
@@ 2.3.3 Mobile defect species
MD simulations going back to the 1990s have established that the general outcome of a dense displacement cascade is a ‘shell’ of SIA clusters surrounding a vacancy-rich core. SIA clusters above a certain size appear as dislocation loops, which in body-centered cubic (BCC) metals are crystallographically ‘perfect’ (unfaulted) and thus potentially glissile along their Burgers vector’s direction wirth2000dislocation, marian2003md. However, in W, SIA clusters are practically immobile for sizes larger than three interstitials faney2014spatially. For their part, vacancies can also appear forming dislocation loops, although their most likely configuration below ¡ is as globular clusters or nanovoids yang2022combined, hu2025new, de2023readdressing, becquart2011modeling, hu2021effect. Similar to their SIA counterparts, vacancy loops are effectively immobile, and, except for select exceptions wei2025resolving, atomistic calculations suggest that only monovacancies possess sufficient mobility among all vacancy clusters oda2014first, wu2024high, long2016first.
In fact, for W there is consistent evidence from isochronal annealing studies that the only species contributing to recovery in stages I, II, and III are, respectively, single SIA, di-interstitials, and monovacancies dicarlo1969stage, KEYS1970260, dausinger1975long, anand1978recovery, seidman1979point, wilson1980situ, nambissan1992positron. TEM experiments in high-purity electron irradiated W have shown that clusters move only in highly driven conditions, which are distinct from conventional ion beam irradiation conditions arakawa2020quantum. Even in atomistic simulations, in pure lattices with no impurities or imperfections, it has been shown that small SIA clusters in W have a reduced mobility compared to other BCC metals zhou2014creeping. Object kinetic Monte Carlo simulations also point to limited or no SIA cluster mobility to find consistency with experimental defect concentration measurements jin2018breaking.
Despite some contradicting work on the diffusion properties of single self-interstitials, in which migration energies as high as 0.3 eV have been reported chen2012stability, zhou2013dynamical, long2017study, wang2021migration, we have found that the migration energies of SIAs that lead to agreement with the present experiments (and are consistent with additional experimental measurements amino2016detection) should be 0.01 eV. One of the main reasons for these low SIA migration energies is that their migration is almost exclusively one-dimensional along directions amino2016detection. It is known that considering rotation as part of the transport mechanism of SIAs leads to larger effective migration energies, as several atomistic calculations have proposed derlet2007multiscale, chen2012stability, long2017study, zhou2024self. However, this is at odds with experimental observations amino2011activation, amino2016detection, heikinheimo2019direct (and other conflicting atomistic studies derlet2007multiscale, becquart2010microstructural, faney2014spatially). Moreover, some authors have suggested that the reason for this low migration energy is that quantum effects may suppress the temperature dependence at low temperatures in W swinburne2017low. With that in mind, the diffusivity parameters for the mobile species in our model, and the sources for their values, are listed in Table 2.
@@
| 0.009 | faney2014spatially, heikinheimo2019direct | ||
| 0.024 | faney2014spatially | ||
| 1.71 | mundy1978self, ventelon2012ab, hossain2014stress, heikinheimo2019direct, maksimenko2022n |
2.3.4 One-dimensional migration of single self-interstitials
As indicated in the previous subsection, SIA mobility is a key property governing the agreement between the experimental measurements and the modeling predictions. With the migration properties firmly established, here we characterize its other crucial aspect: 1D diffusion333Note that in the context of irradiation damage, ‘1D diffusion’ refers to the migration of the defects along rectilinear paths in a 3D setting.. Consideration of one-dimensional migration of self-interstitials implies employing material parameters in the equation shown in Fig. 1 and in eq. (5) that reflect this characteristic of SIA motion. In particular, 1D migration leads to specific coefficients and . In our setting, where the only operational defect sinks are free surfaces bounding the irradiated sample film, the corresponding sink strength, , for SIA migration is borodin1998rate, huang2019rates:
where is the thickness of the irradiated film (approximately 1 mm for the experiments here). The reaction rate between a 1D-moving species (e.g., an SIA) and an immobile object (a vacancy, vacancy cluster, or another SIA cluster) is written as huang2019rates:
where is the diffusivity of (mobile) species ; and are the interaction radii of the two reacting species; and and are their respective concentrations. Note that the above rate closely resembles that of the 3D-3D case for low values of . Following Huang and Marian huang2019rates, the reaction rate between two one-dimensionally migrating species reactions (e.g., single SIAs reacting with one another) is neglected in this work due to their extremely low interaction cross-sections.
@@ 2.3.5 PKA energy distribution tail
The final element in our model that has a strong bearing on the ability of our simulations to capture the experimental size distributions is the need to introduce PKAs with sufficiently high energies to sample the large-size portion of the cluster size distribution. As justified above, due to the limited defect mobility at 290 K in W, the only feasible formation mechanism of large loops at this temperature is through direct injection by displacement cascades. However, as noted by Sand et al. sand2013high and Byggmastar et al. byggmastar2025four, the probability of generating loops with more than 100 defects in W requires PKA energies higher than 100 keV (which is consistent with the value obtained using eq. (2)). From Fig. 2, the probability of sampling energies larger than 100 keV is approximately one in 100,000. Because each ion produces on average 18,373 recoils, roughly ion insertions are needed to produce one 100 keV cascade. Thus, to ensure that the high-energy part of the PKA energy distribution is adequately sampled, a sufficient number of ion insertion events must be simulated (). The number of such events can be estimated from the ion insertion rate, , in the simulation volume defined as:
where is the dpa rate, the atomic density of W, and the simulation volume described in A (all these parameters are given in Table 1). For sufficient statistics, it is desirable to maintain a high ion insertion rate, which, in the context of our simulations, implies using as large a value of as reasonably possible. This imposes numerical demands on the simulations that have to be confronted with appropriate computational resources, but that are essential to adequately sample the large-size defect concentration tails observed experimentally. In this work, we use m-3, i.e., one cubic micron.
@@ 3.1 Cluster size distributions
Figure 6(a) shows the computed SIA and vacancy cluster size distributions at the end of the irradiation period and after post-irradiation annealing for one hour at room temperature. In both cases, the dislocation loop size distributions determined experimentally by the XRDS measurements (corresponding to the ‘after annealing’ simulation results) are shown for comparison. For reference, the figure also shows the extent of TEM-visible cluster size range assuming planar loops, i.e., .
Figure LABEL:sia-scatter displays the SIA cluster size distributions in the ‘as irradiated’ and ‘after annealed’ conditions. The scatter in the model defect size distributions is due to the stochastic nature of our simulations, which, for a given parameter set, yield different but statistically valid results when repeated several times. Furthermore, the statistical variance increases with cluster size due to limited statistics. As the figure shows, the difference between the as-irradiated and annealed size distributions is most visible for cluster sizes with . This is a direct consequence of the high mobility of and defects, whose populations, in the absence of irradiation, are partitioned in the following manner:
@@
Our results show that the modeled defect size distributions match the experimental results, both before and after post-irradiation annealing, for larger cluster sizes ranging from to as large as , where the statistical variance in the modeled distribution becomes significant. Interestingly, the agreement is remarkably accurate if one considers the lower envelope (convex hull) of the simulated statistical distributions. For smaller defects, including the smallest loop size () incorporated into the experimental analysis, the agreement improves after the simulated annealing, owing principally to a decline in the populations of small SIA clusters through dissociation and recombination.
Figure LABEL:vac-scatter shows the vacancy cluster size distributions for the as-irradiated and annealed conditions. Qualitatively, the vacancy cluster population displays a size distribution characteristically similar to its self-interstitial counterpart for larger cluster sizes (). However, consistent with the experimental loop size distributions Schwendeman_2026, the larger vacancy cluster densities predicted by the model are substantially lower than the SIA densities of the same size, and the vacancy size distribution falls off significantly faster with increasing . Thus, the majority of the defects in the TEM-visible size range will be SIA clusters. Another important distinction from the SIA distribution is that the plot shows a stable monovacancy contribution in the as-irradiated condition that reaches very high concentrations. This is a consequence of the negligible mobility of vacancies at 290 K, which leads to a high retention of monovacancies in the material. After the annealing, most of the di-vacancy () population disappears on account of its thermal instability (negative binding energy, see Table 3), and a concentration maximum develops for the cluster. For smaller vacancy cluster sizes (), the model predicts appreciably higher densities with decreasing cluster size than the experiment. This can be attributed to the fact that the XRDS measurements are analyzed to extract the vacancy loop size distribution, specifically (as discussed in Sec. 2.1), while the modeled cluster distribution includes both vacancy loops and globular vacancy clusters.
@@ 3.2 Defect cluster evolution
Figure 6(b) shows the evolution of the SIA and vacancy subpopulations as a function of irradiation time. Each population is partitioned into single point defects ( and ), all clusters (), and prismatic loops larger than the TEM visibility limit (). The evolution of SIA defects is displayed in Fig. LABEL:sia-concentration, where the unstable nature of single interstitials is clear from its relatively low concentrations and highly volatile time evolution. This is, as discussed above, a direct consequence of their high mobility and transient behavior. In contrast, the SIA cluster populations exhibit well-defined, monotonically-increasing trends throughout the entire irradiation process.
The evolution of the vacancy clusters is shown in Fig. LABEL:vac-concentration. In this case, the monovacancy population is seen to be stable and grow at the same rate as larger clusters, although in much higher concentrations. Indeed, more than 90% of vacancies appear and remain as monovacancies in the simulations. The concentration of visible loops is 500 times smaller than that of monovacancies. Clusters of both types are observed to increase linearly, indicating that the system is found in a transient stage of defect accumulation and has not reached saturation yet. A comparison of these results with existing time-resolved experimental data is provided in Sec. 4.3.
It is interesting to note that the fraction of large SIA clusters () increases at a faster rate than the general SIA cluster population (. This is not the case for vacancy clusters, clearly showcasing the increasing importance of SIA loops in acting as effective sinks for single SIAs and small SIA clusters as they progressively accumulate in the material with irradiation dose. This also highlights the time-dependent nature of the defect size distributions, which must be taken into account when interpreting defect cluster concentration evolutions.
@@ 4.1 Capabilities of physics-based numerical simulation model
We have identified four essential factors without which model predictions would not have achieved the high degree of agreement with the experimental data. In no particular order, these factors are: (i) accurately capturing the amount of defects produced in cascades as a function of PKA energy; (ii) a universal scaling law of cluster sizes for defects generated in displacement cascades (independent of PKA energy); (iii) one-dimensional, athermal SIA diffusion; and (iv) consideration of the disproportionate importance of the high-energy tail of the PKA energy distribution function. All four have been implemented in our models, jointly contributing to the accuracy of the simulations and the agreement with experiments. Point (i) ensures that a sufficient number of defects are inserted to account for the large clusters observed experimentally. Point (ii) guarantees that the defect clusters inserted into the simulation volume appear with the correct sizes based on the primary damage mechanisms unveiled by MD studies. Point (iii) maintains a high SI mobility at low temperatures, which controls recombination with vacancies and post-cascade SIA loop growth. Finally, point (iv) plays the crucial role of ensuring that the material is seeded with few but important large immobile dislocation loops, which act as a stable substrate on which further irradiation damage accumulates and evolves. Another factor related to this last point is the importance of considering the full PKA energy distribution, not just the average PKA energy (or any other single representative energy) as has been done in several studies zhang2017radiation, Granberg_2021.
The factors above pertain to the physics of irradiation damage in W, and are in principle agnostic to the model through which they are implemented. These may include okMC becquart2010microstructural, becquart2011modeling, wu2025influence, MFRT Li_Zhou_Ning_Huang_Zeng_Ju_2012, jin2018breaking, li2022towards, or our hybrid method SCD, which builds on advantageous features of both: (i) a mean-field description of the defect populations for computational efficiency and reaching high irradiation doses, and (ii) uses a kinetic Monte Carlo solver to avoid numerical solution bottlenecks in MFRT and provide a meaningful statistical variance to assess numerical errors.
@@
4.2 Comparison between model predictions and XRDS measurements
The simulated SIA and vacancy defect cluster distributions show excellent overall agreement with the XRDS measurements across two orders of magnitude of cluster sizes spanning nearly four orders of magnitude in concentration. For clusters with sizes , the modeling predictions match the experimental results both before and after post-irradiation annealing at 290 K. The distinct SIA and vacancy cluster size distributions measured in this larger size range can therefore be primarily attributed to the differing formation mechanisms of the two cluster types within the cascades. As shown by byggmastar2025four, both clusters types with are produced in cascades initiated by keV PKAs. While Fig. 4(a) shows that the inverse power law size distributions for the two cluster types are very similar, the clustering fraction for vacancies remains substantially lower than that for SIAs in these high energy cascades byggmastar2025four. This crucial distinction is further illustrated by Fig. 6(b) and is ultimately reflected in the differing cluster distribution predictions at the end of the irradiation period. The remarkable agreement between these modeling predictions and the XRDS-measured distributions shown in Fig. 6(a) represents an important experimental validation of the general accuracy of recent cascade physics studies.
For SIA defects (Fig. LABEL:sia-scatter) in the as-irradiated condition, the concentrations of clusters with diverge from the measured distribution with increasingly higher predicted cluster concentrations with decreasing cluster size. This discrepancy is mitigated after post-irradiation annealing at 290 K, suggesting that the combined effect of thermal coarsening and monomer dissociation during a prolonged annealing period substantially affects the concentrations of smaller cluster sizes present in the material over experimental time scales. This has important implications for comparing model predictions with experimental measurements of defect distributions and property evolution—measurements that are often performed ex situ at room temperature, as was the case in this study. The evolution of the vacancy cluster distribution during post-irradiation annealing is more subtle, with Fig. LABEL:vac-scatter showing that the only substantial change in the modeling predictions is the reduction of the concentration of the di-vacancy population by several orders of magnitude due to its thermal instability as discussed above. Instead, the significantly lower clustering fraction for vacancies within the cascades (for all PKAs with energies keV byggmastar2025four) results in a stable and extremely high concentration of monovacancies.
Fig. LABEL:vac-scatter shows that the model also predicts increasingly higher concentrations of vacancy clusters with decreasing size in comparison to the experimental distribution for . This discrepancy is attributed to the analysis of the XRDS measurements extracting the concentrations of vacancy loops, specifically. Importantly, these XRDS measurements do not preclude the presence of small globular vacancy clusters, which have weak distortion fields and therefore do not produce strong XRDS intensities around Bragg reflections. In fact, as discussed by Schwendeman et al. Schwendeman_2026, 85% of the point defects in the XRDS-measured dislocation loop population are found to be SIAs. Assuming roughly equal densities of SIA and vacancy point defects under these irradiation conditions, the large density of ‘missing’ vacancies in the loop population is consistent with PAS measurements (see the following subsection) and the modeling results shown here that indicate a large population of (primarily) monovacancies and small globular clusters is present in the material. Furthermore, by taking the difference between the number of point defects in the XRDS-measured interstitial and vacancy loop populations, a concentration of approximately m-3 vacancy point defects is inferred to be in non-loop structures. This agrees remarkably well with the modeling prediction of approximately m-3 monovancacies after post-irradiation annealing, as illustrated in Fig. LABEL:vac-scatter.
The experimental interstitial and vacancy loop size distributions shown in Fig. 6(a) have been determined from least-squares fits to XRDS measurements using size-dependent cross-section calculations for dislocation loops Schwendeman_dissertation, Schwendeman_2026. Such XRDS cross-section calculations are fundamentally based on detailed atomistic modeling of the structures of defect clusters and the distortions they induce in the surrounding lattice Ehrhart_1982, Schwendeman_dissertation. Like most studies to date, the cluster modeling utilized in the analysis here involves manually constructing the dislocation loops at the center of a lattice and then imposing elastic continuum theory calculations of the defect-induced displacement field on the surrounding atoms Schwendeman_dissertation. However, it is well known that the structure and distortion fields of the smallest defect clusters () can diverge appreciably from this continuum theory based modeling marian2002100. As discussed in detail by Larson Larson_2019 and further by Schwendeman and coworkers Schwendeman_dissertation, Schwendeman_2026, this motivates incorporating MD modeling of the defect clusters into the XRDS cross-section calculations, which can can, in principle, be extended down to point defects and the very smallest cluster sizes with arbitrary configurations. This would enable improved and more detailed comparisons between the modeling and measurements in the smallest cluster size range. Nevertheless, the agreement observed in Fig. 6(a) with the existing experimental analysis, particularly for larger cluster sizes (), is outstanding, which adds a high level of confidence to the underlying physical principles behind the simulation models. Future work will investigate scenarios involving different irradiation conditions –specifically temperature, particle energy, and particle mass– with the ultimate goal of extrapolating these findings to fusion energy conditions in leading conceptual designs.